# Enhancement of quantum sensing in a dissipatively coupled two-mode system

### v1

@@ Abstract
Quantum sensing near exceptional points (EPs) in non-Hermitian systems has shown promising sensitivity enhancements. However, practical applications are often hindered by structural complexity and strict parameter constraints. In this work, we introduce a simplified anti-parity-time (anti-PT) symmetric platform consisting of two independently cavities, which are indirectly coupled to each other by a shared dissipative environment. We demonstrate a significantly enhanced sensing response at the EPs compared to non-EP configurations. This improvement is attributed to the dominant second-order term in the Laurent series expansion of the eigenvalue response to external perturbations—a characteristic feature of higher-order singularities at EPs. This mechanism not only reinforces the foundation for sensitivity enhancement but also offers a structurally compact and robust strategy for quantum sensing. Our results underscore the potential of anti-PT symmetric systems in enabling high-precision sensing technologies and bridging non-Hermitian physics with scalable photonic device platforms.
@@ I INTRODUCTION
Quantum sensing [1, 2, 3] has emerged as a transformative approach in precision measurement, offering unparalleled sensitivity beyond classical limits [4, 5, 6]. Recent advances exploit quantum coherence and entanglement to enable breakthroughs across diverse fields, neutrino mass measurements in particle physics [7], nonlinear spectroscopic techniques via photon-pair correlations [8], optomechanical dark-matter detection in astrophysics, and superconducting circuit-based sensors [9]. Complementary developments in atomic systems (vapor-cell alkali atoms, diamond NV centers, SiC defects [10, 11]) further expand the technological horizon [12, 13] .
@@
Enhancing sensor precision has emerged as a critical research focus in quantum sensing [14, 15, 16]. While Hermitian quantum systems can achieve sensitivity surpassing classical instruments at critical points, recent studies have demonstrated that non-Hermitian systems [17, 18, 19] offer a superior platform for quantum metrology [20, 21, 22, 23]. By leveraging the critical dynamics of gain and loss in open quantum systems [24, 25, 25], these systems exhibit various precision scaling behaviors at critical or exceptional points [26, 27, 28]. However, it is crucial to note that in practical implementations, the inevitable presence of quantum noise in non-Hermitian systems remains a significant factor that cannot be neglected [29, 30]. Nevertheless, despite the dissipation and growing number of noise mitigation strategies [31, 32], the inevitable presence of quantum noise and time constraints still poses significant challenges for quantum sensing. As a paradigmatic non-Hermitian system, parity-time (PT)-symmetric systems [33, 34, 35, 36, 37, 38, 39] have attracted significant attention due to their exceptional points (EPs) [40, 41, 42, 43] -induced eigenvalue degeneracy, which enables exponential sensitivity enhancement and facilitates the design of EP-based quantum sensors [44, 45, 46, 47, 48]. Similarly, anti-PT symmetry [49] has recently emerged as a promising platform for quantum metrology [50, 51]. Recent experimental demonstrations across diverse systems [52, 53, 54, 55] have confirmed the potential of anti-PT-symmetric criticality in achieving enhanced sensing performance [56, 57, 58, 59, 60]. However, the impact of quantum noise and the development of dual-cavity sensing architectures remain underexplored [61].
The advent of dissipative coupling has significantly propelled advancements in quantum sensing [62]. Various devices employing dissipative coupling principles, including optomechanical microdisk-waveguide coupled systems and plasmonic dispersion measurements in coaxial waveguides via single-cavity optical transmission, have demonstrated its efficacy [63, 64]. Furthermore, comprehensive validation across diverse configurations - ranging from cavity-free architectures to single-mode and multimode systems has firmly established the theoretical feasibility of dissipative coupling for quantum-enhanced sensing applications [65].
@@
In this work, we theoretically study the enhanced quantum sensing near the EPs in a dissipatively coupled two-mode system composed of two optical cavities and a waveguide [56, 57]. Unlike conventional PT-symmetric systems that rely on balanced gain and loss, we achieve the anti-PT symmetry by ensuring equal gain or loss in two independent modes through dissipative coupled channels. Using the quantum Langevin equation and the divergence properties of the transformation function, we derive the parameter relations and the probing frequency in the anti-PT symmetric broken/unbroken region. Furthermore, we analyze the quantum Cramér–Rao bound (QCRB) near the probing frequency by using the Laurent expansion approach. We find that when the frequency is close to the probing frequency, the behavior of the QCRB differs significantly across different phases. Such as, the QCRB may exhibit a deeper minimum near the EPs at the probing frequency in the broken phase. However, away from both the EPs and the probing frequency, the QCRB recovers a linear scaling behavior. In the unbroken phase, the QCRB consistently exhibits a linear relationship. In contrast to the above, the QCRB exhibits a quadratic dependence at the EPs due to its second-order nature, which indicates a pronounced enhancement in quantum sensing. This result is consistent with the conventional enhancement of sensing at EPs. Our scheme primarily relies on a shared environment to achieve dissipative coupling between two independent cavities, thereby constructing an anti-PT symmetric system. This proposal constructs the non-Hermitian Hamiltonian without relying on post-selection and is experimentally feasible, with potential implementations in platforms such as superconducting circuits and integrated photonic systems.
@@
The rest of this paper is organized as follows: In Sec. II, we introduce a dissipatively coupled two-mode quantum system, and discuss its anti-PT symmetric properties through the Lindblad master equation framework. In Sec. III, the quantum Fisher information and QCRB are calculated near EPs. Finally, Sec. IV assess the sensitivity near the laser threshold with EPs and we provide a conclusion in Sec. V.
@@ II Theoretical Model and Analysis
The investigated system comprises two spatially separated optical cavities ($a$ and $b$), as illustrated in Fig. 1, where the two cavities are coupled to a shared thermal bath with rate $\Gamma$ [56, 57]. This dissipation architecture induces a dissipative coupled channel alongside the common noise processes. In addition to the respective dissipative environment, each optical cavity mode is also interacted with an independent probing channel with strength $\gamma_{c}$ [66], enabling simultaneous readout and control of the system. The complete time evolution of the system’s density matrix $\rho$ is governed by the following quantum master equation
@@
Figure 1: Two separated optical cavities ($a$ and $b$) interact with each other indirectly via the symmetric dissipation channel. Furthermore, each cavity exhibits their intrinsic loss ($\gamma_{a}$ and $\gamma_{b}$) and the additive probe channel (Channels 11 and 22).
@@
By analogy with the Schrödinger equation, the dynamical equations can be written in compact form $|\dot{\psi}\rangle=-i\mathcal{H}|\psi\rangle$ and $|\psi\rangle=(\langle\hat{a}\rangle,\leavevmode\nobreak\ \langle\hat{b}\rangle)^{\mathrm{T}}$. The effective non-Hermitian Hamiltonian is written as
@@
Figure 2: Eigenvalues $\lambda_{\pm}$. The black solid lines represent the real part of $\lambda_{\pm}$, and the blue dashed lines give the imaginary part of $\lambda_{\pm}$. The points at which the eigenvalues become degenerate are the EPs of the system, i.e., $\delta=\pm 2\Gamma$. The region (orange area) between two EPs corresponds to the unbroken anti-PT symmetric phase.
@@ III Quantum Cramér-Rao Bound and Sensitivity Enhancement Near Exceptional Points
In this dissipatively coupled two-cavity system, the quantum noise originates from multiple channels, such as the input noise through the probe port, the environmental coupling losses, and the noise induced by the dissipative coupling channel. Based on the effective non-Hermitian Hamiltonian $\mathcal{H_{\mathrm{eff}}}$ [Eq. (6)], we derive the quantum dynamical equations for the two cavity modes $a(b)$ via the quantum Langevin equation and the results is written as
$\dot{\hat{a}}$ $=$ $(-i\frac{\delta}{2}-\gamma_{0}-\Gamma)\hat{a}-\Gamma\hat{b}+\sqrt{\gamma_{c}}\hat{a}_{\mathrm{in}}+\sqrt{\gamma_{i}}\hat{N}_{a}+\sqrt{\gamma_{\Gamma}}\hat{N}^{\prime}_{a}$ (10)
$\dot{\hat{b}}$ $=$ $(i\frac{\delta}{2}-\gamma_{0}-\Gamma)\hat{b}-\Gamma\hat{a}+\sqrt{\gamma_{c}}\hat{b}_{\mathrm{in}}+\sqrt{\gamma_{i}}\hat{N}_{b}+\sqrt{\gamma_{\Gamma}}\hat{N}^{\prime}_{b},$ (12)
where $\hat{a}_{\mathrm{in}}(\hat{b}_{\mathrm{in}})$ denotes the input field in each probe channel. $\gamma_{i}=\gamma_{a/b}$ and $\gamma_{\Gamma}$ denote the intrinsic loss of the optical cavity and the dissipative coupling rate, while $\hat{N}_{a/b}$ are the noise operators associated with intrinsic losses, the noise operators $\hat{N}^{\prime}_{a/b}$ arise from fluctuations in the dissipative coupling channel, modeled as vacuum noise.
@@
$\hat{x}_{a}^{\mathrm{in}}=$ $\hat{a}_{\mathrm{in}}+{\hat{a}_{\mathrm{in}}}^{\dagger},$ (17)
$\hat{x}_{b}^{\mathrm{in}}=$ $\hat{b}_{\mathrm{in}}+{\hat{b}_{\mathrm{in}}}^{\dagger},$ (19)
$\hat{y}_{a}^{\mathrm{in}}=$ $-i(\hat{a}_{\mathrm{in}}-{\hat{a}_{\mathrm{in}}}^{\dagger}),$ (21)
$\hat{y}_{b}^{\mathrm{in}}=$ $-i(\hat{b}_{\mathrm{in}}-{\hat{b}_{\mathrm{in}}}^{\dagger}),$ (23)
$\hat{q}_{a/b}=$ $\hat{N}_{a/b}+\hat{N}_{a/b}^{\dagger},$ (25)
$\hat{p}_{a/b}=$ $-i(\hat{N}_{a/b}-\hat{N}_{a/b}^{\dagger}),$ (27)
$\hat{q}^{\prime}_{a/b}=$ $\hat{N}^{\prime}_{a/b}+\hat{N}^{\prime\dagger}_{a/b},$ (29)
$\hat{p}^{\prime}_{a/b}=$ $-i(\hat{N}^{\prime}_{a/b}-\hat{N}^{\prime\dagger}_{a/b}),$ (31)
@@
$V_{\mathrm{i}}=V_{\Gamma}=\begin{bmatrix}1&0&i&0\\ 0&1&0&i\\ -i&0&1&0\\ 0&-i&0&1\end{bmatrix}$. (38)
@@
$\mu_{\mathrm{out}}(\omega)=(I-\gamma_{p}\mathcal{G}_{APT}(\omega))\begin{bmatrix}\hat{x}_{a}^{\mathrm{in}}\\ \hat{x}_{b}^{\mathrm{in}}\\ \hat{y}_{a}^{\mathrm{in}}\\ \hat{y}_{b}^{\mathrm{in}}\end{bmatrix}$. (39)
@@
From Eqs. (37,40), we can know that QFI $I(\omega)$ is only related to the detection input channel $\mu_{\mathrm{in}}(\omega)$ and the transformation function $\mathcal{G}_{APT}(\omega)$. To enforce this divergence, the inverse transformation function $\mathcal{G}^{-1}_{APT}(\omega)$ must become singular at the detection frequency $\omega$. This condition is mathematically expressed through the determinant equation
@@
where $C_{k/m}$ is the expanded coefficient matrix with the pole order $k(m)$. If the transformation function $\mathcal{G}_{APT}(\omega)$ can be expanded to Eq. (55) around the defective point $\omega_{0}$, we can determine that the two parameters in QFI satisfy $V_{\mathrm{out}}^{-1}\sim(\omega-\omega_{0})^{2m}$ and $d\mathcal{G}_{APT}(\omega)/d\omega\sim(\omega-\omega_{0})^{-2m}$ based on Eq. (37). Then we know that the QFI is only related to the order of the transformation function $m$ and exhibits a relationship with $I(\omega)\sim(\omega-\omega_{0})^{-2m}$. We can determine the lower sensitivity limit $\delta\omega$ through the relationship between the QCRB [72, 73, 74] and QFI $\delta\omega=1/\sqrt{I(\omega)}$. Similarly, we can also obtain QCRB that should meet the requirements as

### v2

@@ Abstract
Quantum sensing near exceptional points (EPs) in non-Hermitian systems has shown promising sensitivity enhancements. However, practical applications are often hindered by structural complexity and strict parameter constraints. In this work, we introduce a simplified anti-parity-time (anti-PT) symmetric platform consisting of two independently cavities, which are indirectly coupled to each other by a shared dissipative environment. We demonstrate a significantly enhanced sensing response at the EPs compared to non-EP configurations. This improvement is attributed to the dominant second-order term in the Laurent series expansion of the eigenvalue response to external perturbations- a characteristic feature of higher-order singularities at EPs. This mechanism not only reinforces the foundation for sensitivity enhancement but also offers a structurally compact and robust strategy for quantum sensing. Our results underscore the potential of anti-PT symmetric systems in enabling high-precision sensing technologies and bridging non-Hermitian physics with scalable photonic device platforms.
@@ I INTRODUCTION
Quantum sensing [1, 2, 3] has emerged as a transformative approach in precision measurement, offering unparalleled sensitivity beyond classical limits [4, 5, 6]. Recent advances exploit quantum coherence and entanglement to enable breakthroughs across diverse fields, neutrino mass measurements in particle physics [7], nonlinear spectroscopic techniques via photon-pair correlations [8], optomechanical dark-matter detection in astrophysics, and superconducting circuit-based sensors [9]. Complementary developments in atomic systems, such as vapor-cell alkali atoms, diamond nitrogen-vacancy centers, and defects in silicon carbide, further expand the technological horizon [10, 11, 12, 13].
@@
Enhancing sensor precision has emerged as a critical research focus in quantum sensing [14, 15, 16]. While Hermitian quantum systems can achieve sensitivity surpassing classical instruments at critical points, recent studies have demonstrated that non-Hermitian systems [17, 18, 19] offer a superior platform for quantum metrology [20, 21, 22, 23]. By leveraging the critical dynamics of gain and loss in open quantum systems [24, 25, 25], these systems exhibit various precision scaling behaviors at critical or exceptional points [26, 27, 28]. In non-Hermitian sensors, researchers have conducted the self-consistent quantum analysis of the sensitivity and found that the nonreciprocity could effectively enhance sensing performance [29]. Meanwhile, the influence of noise in practical implementations warrants careful consideration [30, 31]. Despite the growing number of noise mitigation strategies [32, 33], the inevitable presence of quantum noise and time constraints still poses significant challenges for quantum sensing. As a paradigmatic non-Hermitian system, parity-time (PT)-symmetric systems [34, 35, 36, 37, 38, 39, 40] have attracted significant attention due to their exceptional points (EPs) [41, 42, 43, 44] inducing eigenvalue degeneracy, which enables exponential sensitivity enhancement and facilitates the design of EP-based quantum sensors [45, 46, 47, 48, 49]. Similarly, anti-PT symmetry [50] has recently emerged as a promising platform for quantum metrology [51, 52]. Recent experimental demonstrations across diverse systems [53, 54, 55, 56] have confirmed the potential of anti-PT-symmetric criticality in achieving enhanced sensing performance [57, 58, 59, 60, 61]. The advent of dissipative coupling has significantly propelled advancements in quantum sensing [62, 63]. This dissipative coupling mechanism, realized through buses or waveguides [64, 65], enables long-range interactions among multiple modes. However, the impact of quantum noise and the development of dual-mode sensing architectures remain underexplored [66].
@@
In this work, we propose a system where two initially uncoupled modes interact through a dissipative coupling channel, demonstrating its potential for enhanced quantum sensing [57, 58]. Unlike conventional PT-symmetric systems that rely on balanced gain and loss, we achieve the anti-PT symmetry by ensuring equal gain or loss in two independent modes through dissipative coupled channels. Using the quantum Langevin equation and the divergence properties of the transformation function, we derive the parameter relations and the probing frequency in the anti-PT symmetric broken/unbroken region. Furthermore, we analyze the quantum Cramér–Rao bound (QCRB) near the probing frequency by using the Laurent expansion approach. We find that when the frequency is close to the probing frequency, the behavior of the QCRB differs significantly across different phases. Such as, the QCRB may exhibit a deeper minimum near the EPs at the probing frequency in the broken phase. However, away from both the EPs and the probing frequency, the QCRB recovers a linear scaling behavior. In the unbroken phase, the QCRB consistently exhibits a linear relationship. In contrast to the above, the QCRB exhibits a quadratic dependence at the EPs due to its second-order nature, which indicates a pronounced enhancement in quantum sensing. This result is consistent with the conventional enhancement of sensing at EPs. Our scheme primarily relies on a shared environment to achieve dissipative coupling between two independent cavities, thereby constructing an anti-PT symmetric system. This proposal constructs the non-Hermitian Hamiltonian without relying on post-selection and is experimentally feasible, with potential implementations in platforms such as superconducting circuits and integrated photonic systems.
@@
The rest of this paper is organized as follows: In Sec. II, we introduce a dissipatively coupled two-mode quantum system, and discuss its anti-PT symmetric properties through the Lindblad master equation framework. In Sec. III, the quantum Fisher information and QCRB are calculated near EPs. Then, Sec. IV discusses the sensitivity across three distinct regions and includes a supplementary feasibility analysis. Finally, we provide a conclusion in Sec. VI.
@@ II Theoretical Model and Analysis
The studied model comprises two separated subsystems (mode $a$ and $b$), as illustrated in Fig. 1, both coupled to a dissipative channel with a dissipation rate $\Gamma$ [57, 58]. This dissipation architecture induces a dissipative coupled channel alongside the common noise processes. In addition to the respective dissipative environment, each mode is also interacted with an independent probing channel with strength $\gamma_{c}$ [67], enabling simultaneous readout and control of the system. The complete time evolution of the system’s density matrix $\rho$ is governed by the following quantum master equation
@@
Figure 1: Two separated modes ($a$ and $b$) interact with each other indirectly via the symmetric dissipation channel. Furthermore, each mode exhibits their intrinsic loss ($\gamma_{a}$ and $\gamma_{b}$) and the additive probe channel (Channels 11 and 22).
@@
By analogy with the Schrödinger equation, the dynamical equations can be written in compact form $|\dot{\psi}\rangle=-i\mathcal{H}|\psi\rangle$ and $|\psi\rangle=(\langle\hat{a}\rangle,\penalty 10000\ \langle\hat{b}\rangle)^{\mathrm{T}}$. The effective non-Hermitian Hamiltonian is written as
@@
Figure 2: Eigenvalues $\lambda_{\pm}$. The black solid lines represent the real part of $\lambda_{\pm}$, and the blue dashed lines give the imaginary part of $\lambda_{\pm}$. The points at which the eigenvalues become degenerate are the EPs of the system, i.e., $\delta=\pm 2\Gamma$. The region (green area) between two EPs corresponds to the unbroken anti-PT symmetric phase.
@@ III Quantum Cramér-Rao Bound and Sensitivity Enhancement Near Exceptional Points
In this dissipatively coupled two-cavity system, the quantum noise originates from multiple channels, such as the input noise through the probe port, the environmental coupling losses, and the noise induced by the dissipative coupling channel. Based on the effective non-Hermitian Hamiltonian $\mathcal{H_{\mathrm{eff}}}$ Eq. (6), we derive the quantum dynamical equations for the two cavity modes $a(b)$ via the quantum Langevin equation and the result is written as
$\dot{\hat{a}}$ $=$ $(-i\frac{\delta}{2}-\gamma_{0}-\Gamma)\hat{a}-\Gamma\hat{b}+\sqrt{\gamma_{c}}\hat{a}_{\mathrm{in}}+\sqrt{\gamma_{a}}\hat{N}_{a}+\sqrt{\gamma_{\Gamma}}\hat{N}^{\prime}_{a},$ (10)
$\dot{\hat{b}}$ $=$ $(i\frac{\delta}{2}-\gamma_{0}-\Gamma)\hat{b}-\Gamma\hat{a}+\sqrt{\gamma_{c}}\hat{b}_{\mathrm{in}}+\sqrt{\gamma_{b}}\hat{N}_{b}+\sqrt{\gamma_{\Gamma}}\hat{N}^{\prime}_{b},$ (12)
where $\hat{a}_{\mathrm{in}}(\hat{b}_{\mathrm{in}})$ denotes the input field in each probe channel. $\gamma_{i}=\gamma_{a/b}$ and $\gamma_{\Gamma}$ denote the intrinsic loss of each mode and the dissipative coupling rate, while $\hat{N}_{a/b}$ are the noise operators associated with intrinsic losses, the noise operators $\hat{N}^{\prime}_{a/b}$ arise from fluctuations in the dissipative coupling channel, modeled as vacuum noise.
@@
$\hat{x}_{a}^{\mathrm{in}}=$ $\hat{a}_{\mathrm{in}}+{\hat{a}_{\mathrm{in}}}^{\dagger},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \hat{y}_{a}^{\mathrm{in}}=-i(\hat{a}_{\mathrm{in}}-{\hat{a}_{\mathrm{in}}}^{\dagger}),$ (17)
$\hat{x}_{b}^{\mathrm{in}}=$ $\hat{b}_{\mathrm{in}}+{\hat{b}_{\mathrm{in}}}^{\dagger},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \hat{y}_{b}^{\mathrm{in}}=-i(\hat{b}_{\mathrm{in}}-{\hat{b}_{\mathrm{in}}}^{\dagger}),$ (19)
$\hat{q}_{a/b}=$ $\hat{N}_{a/b}+\hat{N}_{a/b}^{\dagger},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \hat{p}_{a/b}=-i(\hat{N}_{a/b}-\hat{N}_{a/b}^{\dagger}),$ (21)
$\hat{q}^{\prime}_{a/b}=$ $\hat{N}^{\prime}_{a/b}+\hat{N}^{\prime\dagger}_{a/b},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \hat{p}^{\prime}_{a/b}=-i(\hat{N}^{\prime}_{a/b}-\hat{N}^{\prime\dagger}_{a/b}),$ (23)
@@
$V_{\mathrm{in}}=V_{\mathrm{i}}=V_{\Gamma}=\begin{bmatrix}1&0&i&0\\ 0&1&0&i\\ -i&0&1&0\\ 0&-i&0&1\end{bmatrix}$. (30)
@@
$\mu_{\mathrm{out}}(\omega)=(I-\gamma_{c}\mathcal{G}_{APT}(\omega))\begin{bmatrix}\hat{x}_{a}^{\mathrm{in}}\\ \hat{x}_{b}^{\mathrm{in}}\\ \hat{y}_{a}^{\mathrm{in}}\\ \hat{y}_{b}^{\mathrm{in}}\end{bmatrix}$. (31)
@@
From Eqs. (29,32), we can know that QFI $I(\omega)$ is only related to the detection input channel $\mu_{\mathrm{in}}(\omega)$ and the transformation function $\mathcal{G}_{APT}(\omega)$. To enforce this divergence, the inverse transformation function $\mathcal{G}^{-1}_{APT}(\omega)$ must become singular at the detection frequency $\omega$. This condition is mathematically expressed through the determinant equation
@@
where $C_{k/m}$ is the expanded coefficient matrix with the pole order $k(m)$. If the transformation function $\mathcal{G}_{APT}(\omega)$ can be expanded to Eq. (47) around the defective point $\omega_{0}$, we can determine that the two parameters in QFI satisfy $V_{\mathrm{out}}^{-1}\sim(\omega-\omega_{0})^{2m}$ and $d\mathcal{G}_{APT}(\omega)/d\omega\sim\mathcal{G}_{APT}^{2}\sim(\omega-\omega_{0})^{-2m}$ based on Eq. (29). Then we know that the QFI is only related to the order of the transformation function $m$ and exhibits a relationship with $I(\omega)\sim(\omega-\omega_{0})^{-2m}$. We can determine the lower sensitivity limit $\delta\omega$ through the relationship between the QCRB [73, 74, 75] and QFI $\delta\omega=1/\sqrt{I(\omega)}$. Similarly, we can also obtain QCRB that should meet the requirements as